im so grateful for all your help! this has helped me in so many ways especially with all the things going on right now. thank you so much !!
(1 point) 35 people are randomly selected and the accuracy of their wristwatches is checked, with positive errors representin 1

(1 point) it is necessary for an automobile producer to estimate the number of miles per gallon achieved by its cars. Suppose2


(1 point) The time needed for college students to complete a certain paper-and-pencil maze follows a normal distribution with3

(1 point) The nicotine content in cigarettes of a certain brand is normally distributed with mean (in milligrams) u and stand4
Answer #1

We would be looking at the first set of questions here on hypothesis test as:

Q1) a) As we are testing here whether the population mean is 0, therefore the test statistic here is computed as:

t^* = \frac{\bar X - \mu_0}{\frac{s}{\sqrt{n}}} = \frac{94 - 0}{\frac{228}{\sqrt{35}}} = 2.4391

Therefore 2.4391 is the required test statistic value here.

b) As we are testing here whether the mean is 0, therefore this is a two tailed test and the p-value here is obtained as:
p = 2P( t34 > 2.4391) = 2*0.01 = 0.02

Therefore 0.02 is the required p-value here.

c) As the p-value here is 0.02 = 0.02 which is the level of significance, therefore the test is significant here and we can reject the null hypothesis here. Therefore we have sufficient evidence here to reject the claim that the mean is equal to 0.